a boat traveled 336 miles downstream and back. the trip downstream took 12 hours. the trip back took 14 hours. what is the speed of the boat in still water?

im clueless???????????

Answer by checkley77(12569)

336/12=28 mph.
336/14=24 mph.
(28-24)/2
4/2=2 mph is the speed of the current.
28-2=26 speed of the boat in still water.
24+2=26 ditto.